The number of pages in a PDF document you create has a discrete uniform distribution from five to nine pages (including the end points). What are the mean and standard deviation of the number of pages in the document ?
1
Expert's answer
2020-06-30T14:37:53-0400

Let X = the random variable denoting the number of pages in the PDF document that has been created


Then we have,


X ~ Uniform{a, b}, a and b are the parameters of the distribution


Now, the expectation and variance of X are given by,


E(X) = "\frac{a+b}{2}"


and Var(X) = "\frac{(b-a+1)^2-1}{12}"


Here we are given,


a = 5, b = 9


Using these values we get,


E(X) = "\frac{5+9}{2}=7"


and Var(X) = "\frac{(9-5+1)^2-1}{12}=2"


Then standard deviation of X = "\sqrt{Var(X)}=\sqrt{2}=1.414" (rounded to 3 decimal places)


Answer: The mean and standard deviation of the number of pages in the document are 7 and 1.414 respectively.